John kicks the ball, and the ball does a projectile motion with an angle of 53o to horizontal with an initial velocity of 10 m/s. What is (a) its maximum height, (b) its maximum distance, and (c) its hangtime?


1
Expert's answer
2021-10-10T15:59:34-0400
"H=\frac{10^2\sin^2{53}}{2(9.8)}=3.25\ m\\
R=\frac{10^2\sin{106}}{(9.8)}=9.81\ m\\
t=2\frac{10\sin{53}}{9.8}=1.63\ s"